OpenStudy (spicytaco1):

What are the solutions to the equation x² = 169? Choose all answers that are correct. A. 13 B. –13 C. 84.5 D. –84.5

OpenStudy (lovelarap):

what do you think?

OpenStudy (calculusisreal):

square root both sides sqrt(x^2)=sqrt(169) x=13

OpenStudy (decentnabeel):

\[x^2=169\] \[\mathrm{For\:}x^2=f\left(a\right)\mathrm{\:the\:solutions\:are\:}x=\sqrt{f\left(a\right)},\:\:-\sqrt{f\left(a\right)}\] \[x=\sqrt{169},\:x=-\sqrt{169}\] \[x=13,\:x=-13\]

OpenStudy (decentnabeel):

@Spicytaco1 are you understand